help! x-y=3
x+y=7
i need help to solve this.

Answer by lucky83(12)

x-y=3
x+y=7
first you need to get either x or y alone in one of the equations.
x=3+y
now you need to substitute this back into the second equation
(3+y)+y=7
2y+3=7
2y=4
y=2
now that you have the y value you can plug it back in to either equation to get the x value.
x+2=7
x=5